OpenStudy (anonymous):

What is the sum? 7/4 + (- 1/5) 31/20 39/20 6/9 -7/20

OpenStudy (anonymous):

31|20

OpenStudy (texaschic101):

7/4 + (-1/5).....when adding/subtracting fractions, you need a common denominator. A number that both denominators go into evenly. That number would be 20. 7/4...now we want the denominator to be 20.....4 times what is 20 ? that would be 5...so multiply 7/4 by 5/5 and you get 35/20. 1/5....so 5 x what is 20 ? that would be 4....so multiply 1/5 by 4/4 and you get 4/20. now set up the problem with your new numbers that have the same denominator ,20. 35/20 + (-4/20) --- pos + neg = neg 35/20 - 4/20 = 31/20 <== your answer

OpenStudy (anonymous):

Thanks! :)

OpenStudy (texaschic101):

no problem :)